ID |
Date |
Author |
Target Material |
Target Oven W.O# |
Source |
Status |
Subject |
Text |
 |
349
|
Tuesday, February 14, 2017, 15:48 |
Laura Lambert | Endcaps | | TARGET BOAT | Done | Sintering TaC coating on Ta endcaps and springs |
|
   |
348
|
Tuesday, February 14, 2017, 15:44 |
Laura Lambert | Endcaps | | TARGET BOAT | Problems resolved | Sintering TaC coating on Ta endcaps and springs |
|
|
347
|
Thursday, February 09, 2017, 23:28 |
John Wong | np-SiC | | TARGET BOAT | Done | Thermotest: SiC-SiO2-Lys and SiC pellets | Thursday, February 09, 2017, 23:26
There are two types of samples being
test for thermal stability: |
|
346
|
Thursday, February 09, 2017, 11:49 |
Laura Lambert | | | | Maintenance | New heatshield with flexible water lines | Thursday, February 09, 2017, 11:47
A new heatshield with flexible water
lines has been put into evap2 last Friday |
|
345
|
Monday, February 06, 2017, 11:17 |
Anders Mjos | | | | Problems resolved | IOC restart after power outagae | Dave Morris did a remote restart
of the IOC after the outage. Turbo pump started. |
|
344
|
Friday, February 03, 2017, 14:18 |
Laura Lambert | Endcaps | | TARGET BOAT | Problems | Sintering TaC coating on Ta endcaps and springs | Friday, February 03, 2017, 14:17
The endcaps are stacked together |
 |
343
|
Tuesday, January 31, 2017, 16:03 |
Laura Lambert | Empty | 39530-10 | HP-FEBIAD | Done | Sintering TaC coated container for backup | Tuesday, January 31, 2017, 13:43
Container coated with TaC using |
 |
342
|
Tuesday, December 20, 2016, 15:04 |
Laura Lambert | | | | Maintenance | RETURN and SUPPLY valves CLOSED for xmas break | Tuesday, December 20, 2016, 15:05
confirmed ~full tank of Ar for glove
box --> ~1600psi |
|
341
|
Friday, December 16, 2016, 14:14 |
Laura Lambert | Ta | 39530-3 | LP-FEBIAD | Done | Conditioning Ta FEBIAD target for shutdown tests | Friday, December 16, 2016, 14:08
Target consists of 470 discs. |
 |
340
|
Thursday, December 01, 2016, 16:08 |
Laura Lambert | Empty | 42218-12 | LP-SIS | Done | sintering TaC coating on UCx#19 container | Thursday, December 01, 2016, 16:06
Container coated with TaC using |
|
339
|
Tuesday, November 22, 2016, 15:58 |
Laura Lambert | Nb | 40282-3 | TARGET BOAT | Done | Conditioning Nb samples for Doug | Tuesday, November 22, 2016, 15:53
Target boat has a layer of niobium |
   |
338
|
Friday, November 18, 2016, 17:28 |
John Wong | np-SiC | | TARGET BOAT | Done | Heat-treatment: Beta-SiC (1um) fiber mats - SAMPLE ID: A1 |
|
|
337
|
Thursday, November 17, 2016, 11:15 |
Laura Lambert | np-SiC | | TARGET BOAT | In progress | Heat-treatment: Beta-SiC (1um) fiber mats - SAMPLE ID: A1 |
|
|
336
|
Tuesday, November 15, 2016, 16:51 |
John Wong | np-SiC | | TARGET BOAT | In progress | Heat-treatment: Beta-SiC (1um) fiber mats - SAMPLE ID: A1 | Tuesday, November 15, 2016, 16:48
Three samples were prepared
(see my logbook#1 page 3, preparing for carbonization). |
|
335
|
Monday, October 24, 2016, 17:52 |
John Wong | np-SiC | | TARGET BOAT | Done | Thermo test of np-SiC raw powder | Monday, October 24, 2016, 17:43
Two samples were prepared and loaded
into the target boat with 1 layer of heat-shied |
 |
334
|
Tuesday, October 18, 2016, 09:23 |
Laura Lambert | Ta | 41790-5 | LP-SIS | Done | Temperature measurements for Ta#50 |
|
|
333
|
Sunday, October 16, 2016, 17:55 |
Aurelia Laxdal | Ta | 41790-5 | LP-SIS | In progress | Temperature measurements for Ta#50 |
|
|
332
|
Saturday, October 15, 2016, 12:42 |
Aurelia Laxdal | Ta | 41790-5 | LP-SIS | In progress | Temperature measurements for Ta#50 |
|
|
331
|
Thursday, October 13, 2016, 18:05 |
Laura Lambert | Ta | 41790-5 | LP-SIS | In progress | Temperature measurements for Ta#50 | Thursday, October 13, 2016, 18:04
Target loaded into evap2
No heat shields. |
|
330
|
Thursday, October 13, 2016, 17:32 |
Laura Lambert | Ta | 41790-5 | LP-SIS | Done | Conditioning Ta#50 |
|
|